OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (–2, 4); Slope: 5 y + 4 = 5x – 10 y + 4 = 5(x + 2) y – 4 = 5(x + 2) y – 4 = 5(x – 2)

OpenStudy (pratyush5):

Equation for point slope form is y - y1 = m(x- x1)

OpenStudy (pratyush5):

Here m = slope and x1, y1 are the point given

OpenStudy (pratyush5):

So you just plugin the values and get the answer

OpenStudy (luigi0210):

Point-slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\] \[\LARGE (x_{1}, y_{1})\] m=slope

OpenStudy (pratyush5):

@Starr_DynastyT Got it ?

OpenStudy (anonymous):

ys i completely understand wht i have gotten is th answer D

OpenStudy (luigi0210):

Try again~

OpenStudy (anonymous):

its incorrect i was surely positive it would be D

OpenStudy (pratyush5):

The negative signs cancel out. Remember that.

OpenStudy (luigi0210):

You ignored your negative~

OpenStudy (pratyush5):